# Sampling-Based Motion Planning with Discrete Configuration-Space Symmetries

### v1

@@ I Introduction
A common property of all these symmetries is that they appear explicitly in the system’s configuration space. In configuration space (or C-space), each coordinate corresponds to an individual degree of freedom, so a single point describes the whole of the positions and orientations of the robot [6]. For many robots, C-space can be interpreted as a bounded subset of Euclidean space. (For example, in the case of a robot arm, where each joint has physical limits to its motion.) But this is not always the case, and the topology of C-space may be inherently non-Euclidean. For example, the C-space of a revolute joint without limits is the unit circle $S^{1}$, and the C-space of a mobile robot is the Lie group $\operatorname{SE}(2)$ (or $\operatorname{SE}(3)$ if the robot can move in 3D space). Finally, closed kinematic chains (or the presence of certain constraints on the motion planning problem) will cause the set of valid configurations to be a curved, measure-zero subset of the full ambient C-space. In all of these cases, the standard strategy is to represent C-space as a smooth manifold [7, 8, 9].
@@ II Related Work
Perhaps the greatest success for leveraging symmetries has been the world of equivariant machine learning, where symmetry-aware learning algorithms can achieve significant improvements in sample complexity, compared to their symmetry-unaware counterparts. These sample complexity gains have been leveraged in grasping [36], simultaneous localization and mapping [37], and point cloud registration [38]. Existing work has also considered the applicability of symmetries in reinforcement and imitation learning for robotics [39], but modifying these strategies to handle object-level equivariances is challenging. Yang et al. [40] used segmentation to find object-centric point clouds, enabling object-level equivariances.
@@ III-A Group Theory
A group is a set $G$ with an associative group law $\cdot:G\times G\to G$. (We often write $gh$ to indicate $g\cdot h$ (for $g,h\in G$), omitting the symbol.) There must be an identity element $e\in G$ and every $g\in G$ must have an inverse. A subgroup is a subset of a group that is itself a group. A group action (on a set $X$) is a binary operation $\alpha:G\times X\to X$ that is compatible with the group law. We say that $G$ acts on $X$ by $\alpha$, and often write $\alpha(g,x)$ as $g\cdot x$ when the action is unambiguous. Note that every group acts on itself via the group law. The orbit of $x\in X$ is the set of points reachable via $G$, denoted $G\cdot x=\left\{g\cdot x:g\in G\right\}$.
We can use groups to describe the symmetries of a rigid body in 2D or 3D. In particular, the configuration of an object in 2D is $\operatorname{SE}(2)$, and we can describe its symmetries as a subgroup of $\operatorname{SO}(2)$ that acts on $\operatorname{SE}(2)$ (and similar with $\operatorname{SO}(3)$ acting on $\operatorname{SE}(3)$ for an object in 3D). Under this action, the orbit of a configuration of the object is the set of all other configurations that are the same under the symmetry.
@@ III-B Symmetry Groups Acting on Manifolds
where $\tilde{x}\in[x]$ is an arbitrary fixed representative. Also, with multiple symmetric objects, we can compute the distance function on a per-object basis. So the cost to compute the distance grows linearly with the number of objects, even though the size of the symmetry group grows quadratically.
@@ IV-A Sampling-Based Planning
    Local sampler: we can draw samples from $B_{r}([q])$ by choosing any $q^{\prime}\in[q]$, sampling $B_{r}(q^{\prime})\subseteq\mathcal{Q}$, and projecting.
@@
Practically speaking, the algorithm will produce a directed graph $G=(V,E)$. Each vertex has a corresponding $q_{v}\in\mathcal{Q}$, the canonical representative of the equivalence class $[q_{v}]$. Note that for each edge in the graph $e=(u,v)$, the minimizing geodesic connecting $q_{u}$ to $q_{v}$ in $\mathcal{Q}$ may not project to the shortest path between $[q_{u}]$ and $[q_{v}]$ in $\mathcal{Q}/G$. So we associate to $e$ the end point $q_{e}\in[q_{v}]$, chosen such that the minimizing geodesic connecting $q_{u}$ to $q_{e}$ projects to the shortest geodesic between $[q_{u}]$ and $[q_{v}]$. This demonstrates why we must internally use a directed graph, as the reverse edge has a different end point.
@@ V Experiments
TABLE II: Performance of symmetry-aware algorithms relative to symmetry-unaware baselines given equal resources (number of samples, connection radii, and neighborhood sizes). For each planner, we show the relative speedup and the relative path length improvement. Higher numbers represent greater speedups and shorter paths, with 1.0 indicating equal performance.
@@
TABLE III: Relative performance of symmetry-aware algorithms when given reduced resources than the symmetry-unaware baselines. Metrics are the same as in Table II.
@@
Finally, to study the scaling with respect to dimension, we consider planning problems with multiple objects with symmetries, for which we must plan paths jointly. We use a bidirectional RRT [52] (capped at 40000 samples). In Table IV, we show the relative improvement in the runtime of the symmetry-aware methodologies, as the dimension of C-space increases. Although the number of samples required still grows exponentially with the dimension, the constant factor reduction in volume of every component still helps the symmetry-aware planners to scale to higher dimensions than the symmetry-unaware baselines. (As discussed in Subsection III-B, the complexity of computing distances grows linearly with the number of copies, even as the order of the symmetry group grows quadratically.)
@@ VI Discussion
In this paper, we have presented a general strategy for adapting sampling-based planning algorithms to handle discrete configuration-space symmetries. We demonstrate that the geometry of such spaces is amenable to efficient planning through the lens of the IMACS framework, as all necessary primitives can be written in closed-form. Rigorous theoretical results demonstrate improvements in sample-complexity, and from a practical perspective, the overall consistency in improvement is very promising.
One clear direction for future work is further theoretical analysis of the performance of sampling-based planners. The geometric properties of the quotient spaces of interest can likely be leveraged for the theoretical analysis of other algorithms and would likely demonstrate similar improvements. Furthermore, the analyses in this paper extend naturally to the kinodynamic setting.
Another direction for future work is extending our results to trajectory optimization, as there are a variety of robotics tasks (especially in contact-rich manipulation) where trajectory optimization has proved more performant than sampling-based planning. Hybrid methods that leverage sampling and trajectory optimization together [3, 4] could naturally be generalized to handle symmetries purely at the roadmap level. And contact-rich planning methods built on Graphs of Convex Sets [53, 54] are suited to modeling symmetries, given the ease with which the symmetry can be reduced to just another discrete decision.
@@
1 3 2 1.121 1.064
2 6 4 1.350 1.099
3 9 8 1.824 1.131
4 12 16 2.573 1.137
5 15 32 2.747 1.101
TABLE IV: Performance of the symmetry-aware bidirectional RRT as the configuration space grows in dimension, with multiple copies of a rectangle. The configuration space is $\operatorname{SE}(2)^{m}$, the the symmetry group is $(C_{2})^{m}$, where $m$ is the number of copies.

### v2

@@ I Introduction
A common property of all these symmetries is that they appear explicitly in the system’s configuration space. In configuration space (or C-space), each coordinate corresponds to an individual degree of freedom, so a single point describes the position and orientation of each part of the robot [6]. For many robots, C-space can be interpreted as a bounded subset of Euclidean space. (For example, in the case of a robot arm, where each joint has physical limits to its motion.) But this is not always the case, and the topology of C-space may be inherently non-Euclidean. For example, the C-space of a revolute joint without limits is the unit circle $S^{1}$, and the C-space of a mobile robot is the Lie group $\operatorname{SE}(2)$ (or $\operatorname{SE}(3)$ if the robot can move in 3D space). Finally, closed kinematic chains (or the presence of certain constraints on the motion planning problem) will cause the set of valid configurations to be a curved, measure-zero subset of the full ambient C-space. In all of these cases, the standard strategy is to represent C-space as a smooth manifold [7, 8, 9].
@@ II Related Work
Perhaps the greatest success for leveraging symmetries has been the world of equivariant machine learning, where symmetry-aware learning algorithms can achieve significant improvements in sample complexity, compared to their symmetry-unaware counterparts. These sample complexity gains have been leveraged in grasping [36], simultaneous localization and mapping [37], and point cloud registration [38]. Existing work has also considered the applicability of symmetries in reinforcement and imitation learning for robotics [39], but modifying these strategies to handle object-level equivariances is challenging. Yang et al. [40] used segmentation to find object-centric point clouds, enabling object-level equivariances, but this approach can only work for a single symmetric object.
@@ III-A Group Theory
A group is a set $G$ with an associative group law $\cdot:G\times G\to G$. (We often write $gh$ to indicate $g\cdot h$ (for $g,h\in G$), omitting the symbol.) There must be an identity element $e\in G$ and every $g\in G$ must have an inverse. We use $|G|$ to denote the cardinality, or order, of $G$. A subgroup is a subset of a group that is itself a group. A group action (on a set $X$) is a binary operation $\alpha:G\times X\to X$ that is compatible with the group law. We say that $G$ acts on $X$ by $\alpha$, and often write $\alpha(g,x)$ as $g\cdot x$ when the action is unambiguous. Note that every group acts on itself via the group law. The orbit of $x\in X$ is the set of points reachable via $G$, denoted $G\cdot x=\left\{g\cdot x:g\in G\right\}$.
We can use groups to describe the symmetries of a rigid body in 2D or 3D. In particular, the configuration of an object in 2D is described by the group $\operatorname{SE}(2)$, and the subgroup $\operatorname{SO}(2)\subseteq\operatorname{SE}(2)$ is the set of orientations. We can describe the object’s symmetries as a subgroup of $\operatorname{SO}(2)$ that acts on $\operatorname{SE}(2)$ (respectively, $\operatorname{SO}(3)$ and $\operatorname{SE}(3)$ for an object in 3D). Under this action, the orbit of a configuration of the object is the set of all other configurations that are the same under the symmetry.
@@ III-B Symmetry Groups Acting on Manifolds
where $\tilde{x}\in[x]$ is an arbitrary fixed representative. Also, with multiple symmetric objects, we can compute the distance function on a per-object basis. So the cost to compute the distance grows linearly with the number of objects, even though the size of the symmetry group grows exponentially.
@@ IV-A Sampling-Based Planning
    Local sampler: we can draw samples from the ball of radius $r$ about $q$, $B_{r}([q])$, by choosing any $q^{\prime}\in[q]$, sampling $B_{r}(q^{\prime})\subseteq\mathcal{Q}$, and projecting.
@@
Practically speaking, the algorithm will produce a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$. Each vertex has a corresponding $q_{v}\in\mathcal{Q}$, the canonical representative of the equivalence class $[q_{v}]$. Note that for each edge in the graph $e=(u,v)$, the minimizing geodesic connecting $q_{u}$ to $q_{v}$ in $\mathcal{Q}$ may not project to the shortest path between $[q_{u}]$ and $[q_{v}]$ in $\mathcal{Q}/G$. So we associate to $e$ the end point $q_{e}\in[q_{v}]$, chosen such that the minimizing geodesic connecting $q_{u}$ to $q_{e}$ projects to the shortest geodesic between $[q_{u}]$ and $[q_{v}]$. This demonstrates why we must internally use a directed graph, as the reverse edge has a different end point.
@@ V Experiments
TABLE II: Performance of symmetry-aware algorithms relative to symmetry-unaware baselines given equal resources. For each planner, we show the relative speedup and the relative path length improvement. Higher numbers represent greater speedups and shorter paths, with 1.0 indicating equal performance. Due to space constraints, additional experiments and complete results (including Radius-PRM*) are available online at https://cohnt.github.io/projects/symmetries.html.
@@
TABLE III: Relative performance of symmetry-aware algorithms when given reduced resources than the symmetry-unaware baselines. Metrics are the same as in Table II. (RRT is omitted as no strategy for budgeting its available resources is presented.)
@@
Finally, to study the scaling with respect to dimension, we consider planning problems with multiple objects with symmetries, for which we must plan paths jointly. We use a bidirectional RRT (BiRRT) [52] (capped at 40000 samples). In Table IV, we show the relative improvement in the runtime of the symmetry-aware methodologies, as the dimension of C-space increases. Although the number of samples required still grows exponentially with the dimension, the constant factor reduction in volume of every component still helps the symmetry-aware planners to scale to higher dimensions than the symmetry-unaware baselines. (As discussed in Subsection III-B, the complexity of computing distances grows linearly with the number of copies, even as the order of the symmetry group grows exponentially.)
@@ VI Discussion
In this paper, we have presented a general strategy for adapting sampling-based planning algorithms to handle discrete configuration-space symmetries. We demonstrate that the geometry of such spaces is amenable to efficient planning through the lens of the IMACS framework, as all necessary primitives can be written in closed-form. Rigorous theoretical results demonstrate improvements in sample-complexity, and comprehensive experimental results verify the theory leads to dramatically improved runtimes for the popular RRT, KNN-PRM*, and Radius-PRM* planning algorithms.
One direction for future work is studying the performance of other sampling-based planners with symmetries. Our analysis is mostly based on properties of the quotient space, so these other approaches should see similar improvements, even in the kinodynamic setting.
Another direction for future work is extending our results to trajectory optimization, as there are a variety of robotics tasks (especially in contact-rich manipulation) where such methods have proved more performant than sampling-based planning. Hybrid methods that leverage sampling and trajectory optimization together [3, 4] could naturally be generalized to handle symmetries purely at the roadmap level. And contact-rich planning methods built on Graphs of Convex Sets [53, 54] are suited to modeling symmetries, given the ease with which the symmetry can be reduced to just another discrete decision.
@@
1 3 2 1.121x 1.064x
2 6 4 1.350x 1.099x
3 9 8 1.824x 1.131x
4 12 16 2.573x 1.137x
5 15 32 2.747x 1.101x
TABLE IV: Performance of symmetry-aware BiRRT as the configuration space grows in dimension, with multiple copies of a rectangle. The configuration space is $\operatorname{SE}(2)^{m}$, the the symmetry group is $(C_{2})^{m}$, where $m$ is the number of copies.